OpenStudy (anonymous):

3 tons per day = ??? ounces per second

OpenStudy (xishem):

We can use dimensional analysis:\[3\frac{tons}{day}\times (\frac{1day}{24hr})\times(\frac{1hr}{3600s})\times (\frac{2000lb}{1ton})\times (\frac{16oz}{1lb})=1.11\frac{oz}{s}\]You can see from the dimensional analysis that all the units cancel out except ounces on top and seconds in the basement.